Contract the expressions. That is, use the properties of logarithms to write each expression as a single logarithm with a coefficient of 1
ln 3 − 2(ln 4 + ln 8)

Answer by oscargut(2103)

ln 3 − 2(ln 4 + ln 8)
= ln(3) − 2(ln(4*8))
= ln(3) -2ln(32)
= ln(3) -ln(32^2)
= ln(3) -ln(1024)
= ln (3/1024)
Answer: ln (3/1024)
You can ask me more at: mthman@gmail.com
Thanks